# Galton-Watson processes, simple varieties of trees and Khinchin families

### v3

@@ Abstract.
In this note, we introduce a unified analytic framework that connects Bienaymé–Galton–Watson processes, simple varieties of trees and Khinchin families. Using Lagrange’s inversion formula, we derive new coefficient-based expressions for extinction probabilities and reinterpret them as boundary phenomena tied to the domain of the inverse of the solution to Lagrange’s equation. This perspective reveals a structural bridge between combinatorial and probabilistic models, simplifying classical arguments and yielding new results. It also leads to a computationally efficient method for simulating Galton–Watson processes via power series coefficients.
@@ 1. Introduction
To each power series $f\in\mathcal{K}$ (see the definition below) with radius of convergence $R>0$ we associate a family of random variables $(X_{t})_{t\in[0,R)}$, the Khinchin family associated to $f$. For each $t\in(0,R)$ the random variable $X_{t}$ has mass function
$\mathbf{P}(X_{t}=n)=\frac{a_{n}t^{n}}{f(t)}\,,\quad\text{ for any}n\geq 1\,.$
with $X_{0}\equiv 0$. The theory of Khinchin families examines the behaviour of this family of random variables, and its normalized version, as the parameter $t\uparrow R$. See, for instance, [3], [4] and [5].
@@
here $A_{n}$ is the $n$-th coefficient of the solution of Lagrange’s equation $g(z)=z\psi(g(z))$ with data $\psi\in\mathcal{K}$. These coefficients can be computed in terms of $\psi$ by using Lagrange’s inversion formula. We refer the reader to Theorem 2 below for a detailed statement and a proof of this result.
For each fixed $\psi\in\mathcal{K}$, the coefficients $A_{n}$ encode a simple variety of trees, that is, the sum of certain fixed weights over all rooted plane trees with $n$ nodes, see Proposition A below. Specific choices of $\psi\in\mathcal{K}$ yield concrete families of combinatorial trees; for instance, $\psi(z)=e^{z}$ corresponds to rooted labeled Cayley trees, while $\psi(z)={1}/{(1-z)}$ corresponds to rooted plane trees.
It turns out that the radius of convergence of the solution to Lagrange’s equation with data $\psi$ is intimately related to the extinction probability of the corresponding Galton–Watson process with offspring distribution $Y_{t}$, marking the critical parameter at which the process shifts from almost sure extinction to positive survival probability. This phenomenon bears a clear resemblance to phase transitions in statistical mechanical models.
@@
Here, we simultaneously consider all possible Galton–Watson processes with offspring distribution given by a Poisson random variable with parameter $t>0$. Notice that $q(t)\equiv 1$, for any $t\leq\tau=1$ and also that the radius of convergence of $g(z)$ is given by $\rho=\tau/\psi(\tau)=1/e$, see Theorem 1 below. When we cross the radius of convergence of $g(z)$, that is, when $t>\tau=1$, the inversion relation ${t}/{\psi(t)}=g^{-1}(t)$ no longer holds, and consequently the extinction probability should satisfy the inequality $q(t)<1$. Observe that for $t>\tau=1$, the function $t/\psi(t)$ is strictly decreasing, see Lemma 4.1 below.
After deriving formula (1.1), we analyze the probability of extinction function in detail, establishing its continuity, differentiability (except at one point) and characterizing its asymptotic behavior as the parameter $t\uparrow R_{\psi}$. An explicit upper bound on the extinction probability function is also derived. Moreover, we establish several precise relationships between Khinchin families and random variables arising in Galton–Watson processes. We identify some of these random variables as appropriately reparameterized Khinchin families.
Beyond its theoretical interest, this framework offers an efficient method for simulating Galton-Watson processes by leveraging power series and the numerical computation of coefficients. We believe that this perspective will be valuable to both combinatorialists and probabilists, as it provides an explicit bridge between these viewpoints.
This note is self-contained and intended to be accessible to researchers with backgrounds in probability, combinatorics, or complex analysis. Section 2 introduces the fundamental concepts of Khinchin families. Section 3 reviews rooted plane trees and Galton–Watson processes. Section 4 recalls Lagrange’s inversion formula and gives a closed form expression for the radius of convergence of the power series that solves Lagrange’s equation. Section 5 provides a combinatorial interpretation of the solution to Lagrange’s equation in terms of weighted rooted plane trees. Finally, Section 6 constitutes the core of this note: it develops the parametric Galton–Watson process, derives explicit coefficient-based formulas for various probabilities, including extinction probabilities, and explores their analytic and probabilistic behaviour.
@@ 2. Khinchin families
In this section, we introduce the key elements of the theory of Khinchin families that are necessary to developed this note. A more comprehensive treatment can be found in [3], [4], [5] and [8] and more recently, in [12] and [13].
@@ 3.3. Galton–Watson processes
A Galton–Watson process begins with a single ancestor whose offspring count $Y$ follows the distribution $(p_{k})_{k\geq 0}$; each descendant independently reproduces according to the same law in successive generations, see, for instance, [2]. If $Z_{n}$ denotes the population at generation $n$, then
@@
with $Y_{j}$ i.i.d. copies of $Y$. Its mean $m=\mathbf{E}[Y]$ determines extinction: if $m\leq 1$, extinction occurs almost surely ($q=1$); if $m>1$, the extinction probability $q<1$. The probability $q$ is the minimal solution of the equation
@@
For further details about these processes see, for instance, [2].
@@ 4. Lagrange’s inversion formula
Then, the coefficients of $g(z)$ satisfy the recurrence:
@@ 4.1. Function $t/\psi(t)$
Lagrange’s equation gives that $g$ and $z/\psi(z)$ are inverse of each other and thus that
$g(z/\psi(z))=z\,,\quad\mbox{for $|z|\leq\tau$}\,.$
In particular,
$g(t/\psi(t))=t\,,\quad\mbox{for $t\in[0,\tau]$}\,.$
The image of the interval $[0,\rho]$ by $g$ is the interval $g([0,\tau/\psi(\tau)])=[0,\tau]$.
@@ 4.2. Radius of convergence of $g$
Let $\psi\in\mathcal{K}$ be a power series with radius of convergence $R_{\psi}>0$. Denote $g$ the solution of Lagrange’s equation with data $\psi$, that is,
$g(z)=z\psi(g(z))\,,$
then the radius of convergence of $g$ is given by
@@
More generally: for power series $\psi\in\mathcal{K}^{\star}$, if $\rho=\tau/\psi(\tau)\geq 1$, then $b_{1}\leq 1$. This is equivalent to the claim: $b_{1}\geq 1$ implies $\rho\leq 1$.
@@ 4.3. Otter-Meir-Moon asymptotic formula
so it is natural to set $g(\rho)=\tau$. We will use this relation repeteadly along this note. Observe that here we use that $g$ has non-negative Taylor coefficients.
@@ 5. Lagrange’s equation and simple varieties of trees
Alternatively some authors define these simple varieties for power series $\psi\in\mathcal{K}^{\star}$, this is not the case here.
@@
$\omega(a)=\prod_{j=0}^{\infty}b_{j}^{k_{j}(a)}\,,\quad\text{ for any }a\in\mathcal{G}\,.$
@@ 6.2. Probability generating functions
$g_{t}(z)=\frac{g(tz/\psi(t))}{t}\,,\quad\text{ for all }t\in[0,R_{\psi})\text{ and }|z|\leq 1\,.$
@@ 9. Conclusion and further directions
The ideas developed in this note extend naturally to subclasses of rooted plane trees defined by various types of restrictions (and also to random forests). Moreover, when combined with the results in [8], this perspective yields precise asymptotic formulas across a broad range of settings. A key strength of the approach is that it unifies and simplifies classical arguments, often reducing them to algebraic manipulations, while making the connection between combinatorial and probabilistic viewpoints explicit. The alternative proof of the Otter–Meir–Moon Theorem in [8], based on Khinchin families and including several extensions, serves as a clear illustration of this principle. The same methodology can be applied to derive asymptotic formulas for the probabilities associated to Galton-Watson processes studied here, and also to simulate them efficiently using power series coefficients.
Finally, we believe that the underlying structure revealed by these results is broadly applicable, suggesting that this methodology may be extended to other weighted combinatorial classes through the framework of Khinchin families. For further details about the theory of Khinchin families see [3], [4], [5], [8], [12] and [13].

### v5

@@ Abstract.
In this note, we introduce a unified analytic framework that connects Bienaymé–Galton–Watson processes, simple varieties of trees and Khinchin families. Using Lagrange’s inversion formula, we derive new coefficient-based expressions for extinction probabilities and reinterpret them as boundary phenomena tied to the domain of the inverse of the solution to Lagrange’s equation. This perspective reveals an additional link between combinatorial and probabilistic models, simplifying classical arguments and yielding new results. It also leads to a computationally efficient method for simulating Galton–Watson processes via power series coefficients.
@@ 1. Introduction
A Bienaymé-Galton–Watson (henceforth Galton–Watson) process is a branching model where every individual independently produces a random number of children following a given offspring distribution. These processes can be traced back to Bienaymé’s work on extinction probabilities in the 1840s and were further developed by F. Galton and H. W. Watson in 1874–1875, see [3, 11, 12] and [27]. For a thorough modern treatment of branching processes, see [2]; for further historical background, see [14, 15].
Their motivating question can be summarized as: Will a given family line eventually die out?. It is a classical result that the extinction probability $q$ of a Galton-Watson process is the smallest non-negative fixed point of the offspring’s probability‐generating function $G_{Y}(z)$, namely
$G_{Y}(q)=\sum_{n=0}^{\infty}\mathbf{P}(Y=n)\,q^{n}=q.$
The first complete proof of this result was given by J. F. Steffensen in his 1930 note [25] (in Danish) and its 1933 expansion [26], see also the unpublished proof by C. M. Christensen in [14]. For additional historical details and an English translation of Steffensen’s 1930 note, see Guttorp’s report [14].
To each power series $f\in\mathcal{K}$ (see the definition below) with radius of convergence $R>0$ we can associate a family of random variables $(X_{t})_{t\in[0,R)}$, the Khinchin family associated to $f$. For each $t\in(0,R)$ the random variable $X_{t}$ has mass function
$\mathbf{P}(X_{t}=n)=\frac{a_{n}t^{n}}{f(t)}\,,\quad\text{ for any}n\geq 0\,,$
with $X_{0}\equiv 0$. The theory of Khinchin families examines the behaviour of this family of random variables, and its normalized version, as the parameter $t\uparrow R$. See, for instance, [4, 5, 6, 9], and more recently [17, 18].
@@
here $A_{n}$ is the $n$-th coefficient of the solution of Lagrange’s equation $g(z)=z\psi(g(z))$ with data $\psi\in\mathcal{K}$. These coefficients can be computed in terms of $\psi$ by using Lagrange’s inversion formula. We refer the reader to Theorem 2 below for a detailed statement and proof of this result.
A simple variety of trees is a family of weighted rooted plane trees in which each node receives a weight that depends on its number of children (outdegree). The weight of the whole tree is the product of the different node-weights, see [10, p. 452]. Throughout this note we choose those node-weights to be the coefficients of a power series $\psi\in\mathcal{K}$ (see eq. (5.1) below). For such a fixed $\psi$, the sequence $A_{n}$ records the sum of the weights over all rooted plane trees with $n$ nodes (see Proposition A below). Specific choices of $\psi$ yield concrete families of combinatorial trees; for instance, $\psi(z)=e^{z}$ correspond to rooted labelled Cayley trees, while $\psi(z)=1/(1-z)$ correspond to rooted plane trees.
It turns out that the radius of convergence of the solution to Lagrange’s equation with data $\psi$ is intimately related to the extinction probability of the corresponding Galton–Watson process with offspring distribution $Y_{t}$, marking the critical parameter at which the process shifts (phase transition) from almost sure extinction to positive survival probability.
@@
Here, we simultaneously consider all possible Galton–Watson processes with offspring distribution given by a Poisson random variable with parameter $t>0$. Notice that $q(t)\equiv 1$, for any $t\leq\tau=1$ (to wit, the series above is actually identical to 1) and also that the radius of convergence of $g(z)$ is given by $\rho=\tau/\psi(\tau)=1/e$, see Theorem 1 below. When we cross the radius of convergence of $g(z)$, that is, when $t>\tau=1$, the inversion relation ${t}/{\psi(t)}=g^{-1}(t)$ no longer holds, and consequently the extinction probability should satisfy the inequality $q(t)<1$. Observe that for $t>\tau=1$, the function $t/\psi(t)=te^{-t}$ is strictly decreasing, see Lemma 4.1 below.
After deriving formula (1.1), we analyze the probability of extinction function in detail, establishing its continuity, differentiability (except at one point) and characterizing its asymptotic behavior as the parameter $t\uparrow R_{\psi}$. An explicit upper bound on the extinction probability function is also derived. Moreover, we establish several precise relationships between Khinchin families and random variables arising in Galton–Watson processes. We identify some of these random variables as appropriately reparameterized Khinchin families. Beyond its theoretical interest, this framework offers an efficient method for simulating Galton-Watson processes by means of power series and the numerical computation of its coefficients. We believe that this perspective will be valuable to both combinatorialists and probabilists, as it provides an explicit bridge between these viewpoints.
This note aims to be self-contained and accessible to researchers with backgrounds in probability, combinatorics, or complex analysis. Section 2 introduces the fundamental concepts of Khinchin families. Section 3 reviews rooted plane trees and Galton–Watson processes. Section 4 recalls Lagrange’s inversion formula and gives a closed form expression for the radius of convergence of the power series that solves Lagrange’s equation. Section 5 provides a combinatorial interpretation of the solution to Lagrange’s equation in terms of weighted rooted plane trees. Finally, Section 6 constitutes the core of this note: it develops the parametric Galton–Watson process, derives explicit coefficient-based formulas for various probabilities, including extinction probabilities, and explores their analytic and probabilistic behaviour.
@@ 2. Khinchin families
In this section, we introduce the key elements of the theory of Khinchin families that are necessary to develop this note. A more comprehensive treatment can be found in [4], [5], [6] and [9] and more recently, in [17] and [18].
@@ 3.3. Galton–Watson processes
A Galton–Watson process begins with a single ancestor whose offspring count $Y$ follows the distribution $(p_{k})_{k\geq 0}$; each descendant independently reproduces according to the same law in successive generations, see, for instance, [2] and [21]. If $Z_{n}$ denotes the population at generation $n$, then
@@
with $Y_{j}$ i.i.d. copies of $Y$ and $Z_{0}=Y$. The process $(Z_{n})_{n\geq 0}$ is a Galton-Watson process. The mean of the offspring distribution $m=\mathbf{E}[Y]$ determines extinction: if $m\leq 1$, extinction occurs almost surely ($q=1$); if $m>1$, the extinction probability $q<1$. The probability $q$ is the minimal solution of the equation
@@
Every realization of the process corresponds to a rooted plane tree that encodes its parent–child relations. For further details about these processes see, for instance, [2] and [21].
@@ 4. Lagrange’s inversion formula
Then, the coefficients of $g(z)$ satisfy the relation:
@@ 4.1. Function $t/\psi(t)$
Lagrange’s equation gives that $g$ and $z/\psi(z)$ are inverse of each other and thus, for power series $\psi\in\mathcal{K}^{\star}$, we have that
$g(z/\psi(z))=z\,,\quad\mbox{ for }z\in\Omega_{g}=g(\mathbb{D}(0,\rho))\subseteq\mathbb{D}(0,\tau)\,.$
Invoking Lagrange’s equation together with the fact that $\psi\in\mathcal{\mathcal{K}}$, one sees that $\psi$ cannot vanish in $\Omega_{g}$. In fact, since $g$ has non-negative coefficients, we conclude that
$g(t/\psi(t))=t\,,\quad\mbox{for $t\in[0,\tau]$}\,,$
that is, the image of the interval $[0,\rho]$ by $g$ is the interval $g([0,\tau/\psi(\tau)])=[0,\tau]$. One may verify this claim directly (using that $g$ is increasing and continuous on the closure of $\mathbb{D}(0,\rho)$), or else deduce it as a corollary of Theorem B below.
@@ 4.2. Radius of convergence of $g$
Let $\psi\in\mathcal{K}$ be a power series with radius of convergence $R_{\psi}>0$. Denote $g$ the solution of Lagrange’s equation with data $\psi$, that is, $g(z)=z\psi(g(z))$ , then the radius of convergence of $g$ is given by
@@
More generally: for power series $\psi\in\mathcal{K}^{\star}$ such that $b_{1}\geq 1$ we have that $\rho\leq 1$.
@@ 4.3. Otter-Meir-Moon asymptotic formula
so it is natural to set $g(\rho)=\tau$. We will use this relation repeteadly along this note.
@@ 5. Lagrange’s equation and simple varieties of trees
Alternatively some authors define these simple varieties for power series $\psi\in\mathcal{K}^{\star}$, this is not the case here. For further details see, for instance, [8], [10, p. 452] and [16].
@@
(5.1) $\omega(a)=\prod_{j=0}^{\infty}b_{j}^{k_{j}(a)}\,,\quad\text{ for any }a\in\mathcal{G}\,.$
@@ 6.2. Probability generating functions
$g_{t}(z)=\frac{g(tz/\psi(t))}{t}\,,\quad\text{ for all }t\in(0,R_{\psi})\text{ and }|z|\leq 1\,.$
@@ 9. Conclusion and further directions
The ideas developed in this note extend naturally to subclasses of rooted plane trees defined by various types of restrictions (and also to random forests). Moreover, when combined with the results in [9], this perspective yields precise asymptotic formulas across a broad range of settings. A key strength of the approach is that it unifies and simplifies classical arguments, often reducing them to algebraic manipulations, while making the connection between combinatorial and probabilistic viewpoints explicit. The alternative proof of the Otter–Meir–Moon Theorem in [9], based on Khinchin families and including several extensions, serves as a clear illustration of this principle. The same methodology can be applied to derive asymptotic formulas for the probabilities associated to Galton-Watson processes studied here, and also to simulate their probabilities efficiently using power series coefficients.
Finally, we believe that the underlying structure revealed by these results is broadly applicable, suggesting that this methodology may be extended to other weighted combinatorial classes through the framework of Khinchin families. For further details about the Theory of Khinchin families see [4, 5, 6, 9, 17, 18].